A new cruise ship line has just launched 3 new ships the pacific paradise, the caribbean paradise, and the mediterranean paradise. the caribbean paradise has 13 more deluxe staterooms than the pacific paradise. the mediterranean paradise has 31 fewer deluxe staterooms than four times the number of deluxe staterooms on the pacific paradise. find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 606.
1
Expert's answer
2012-07-24T11:22:29-0400
Let X is the number of deluxe staterooms on the pacific paradise, Y -
on the
caribbean paradise, and Z - on the mediterranean paradise.
Y = X + 13
Z =
4*X - 31
X + Y + Z = 606 <=> X + (X + 13) + 4*X - 31 = 606 <=>
6*X - 18 = 606
<=> 6*X = 624 <=> X = 624 / 6 = 104.
Y = 104 +
13 = 117, Z = 4*104 - 31 = 385.
Answer: 104; 117; 385.